# ChatQA 2: Bridging the Gap to Proprietary LLMs in Long Context and RAG Capabilities

### v1

@@ Abstract
In this work, we introduce ChatQA 2, a Llama3-based model designed to bridge the gap between open-access LLMs and leading proprietary models (e.g., GPT-4-Turbo) in long-context understanding and retrieval-augmented generation (RAG) capabilities. These two capabilities are essential for LLMs to process large volumes of information that cannot fit into a single prompt and are complementary to each other, depending on the downstream tasks and computational budgets. We present a detailed continued training recipe to extend the context window of Llama3-70B-base from 8K to 128K tokens, along with a three-stage instruction tuning process to enhance the model’s instruction-following, RAG performance, and long-context understanding capabilities. Our results demonstrate that the $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ model achieves accuracy comparable to GPT-4-Turbo-2024-0409 on many long-context understanding tasks and surpasses it on the RAG benchmark. Interestingly, we find that the long-context retriever can alleviate the top-k context fragmentation issue in RAG, further improving RAG-based results for long-context understanding tasks. We also provide extensive comparisons between RAG and long-context solutions using state-of-the-art long-context LLMs.
@@ 1 Introduction
In recent developments, the trend of extending the context window length in LLMs has gained remarkable traction within both the industrial and research communities. All leading proprietary LLMs support very large context window, allowing them to accommodate several hundred pages of text in a single prompt. For example, GPT-4 Turbo (OpenAI, 2023) and Claude 3.5 Sonnet offer a 128K and 200K context window, respectively. Meanwhile, Gemini 1.5 Pro (Gemini-Team, 2024) impressively supports up to a 10M context. Open-access LLMs have also made significant strides to keep up (01.AI et al., 2024; Alibaba-QWen, 2024). For instance, QWen2-72B-Instruct (Alibaba-QWen, 2024) and Yi-34B (01.AI et al., 2024) support 128K and 200K context windows, respectively. However, the training data and technical details for these models are missing, making reproduction challenging. In addition, these models have mostly been evaluated on synthetic tasks, like Needle in a Haystack (Kamradt, 2023) test, which does not accurately represent real-world downstream task performance. For example, previous studies shows a noticeable gap between open-access LLMs and leading proprietary models on real-world long context understanding tasks (Zhang et al., 2024a; Hsieh et al., 2024). In this work, we focus on bridging the gap between the open-access Llama-3 and proprietary GPT-4 Turbo on real-world long context understanding tasks.
@@
The long-context capability of LLMs is sometimes considered a rival technique to retrieval-augmented generation (RAG). However, from a pragmatic perspective, these techniques complement each other. An LLM with a long context window can either process large volumes of text as a prompt or utilize retrieval methods to efficiently extract relevant information from the extensive text, depending on the downstream tasks and accuracy vs. efficiency trade-offs. RAG has efficiency advantages and can easily retrieve relevant contexts for query-based tasks (e.g, QA) from billions of tokens, a feat that long context models cannot achieve. Meanwhile, long context models are good at tasks such as summarizing entire documents, where RAG may not perform as well. As a result, the state-of-the-art LLM needs to excel at both capabilities, providing options for different downstream tasks based on accuracy and efficiency requirements. In a previous study, the open-source ChatQA 1.5 (Liu et al., 2024) model can surpass GPT-4-Turbo on RAG tasks. In this work, we present ChatQA 2, which possesses both GPT-4-Turbo level long context understanding capability and RAG performance.
Xu et al. (2024) extended the context window of Llama2 (Touvron et al., 2023b) to 16K and 32K tokens, and studied the interplay between RAG and long-context LLMs. It demonstrates that the RAG method can improve the generation accuracy and inference efficiency of a GPT-3.5-turbo-16k level long context model for QA and query-based summarization tasks. In this work, we extend this study by pushing the long context LLM to GPT-4-Turbo level with 128K context window and combining it with a state-of-the-art long-context retriever for RAG.
@@
    We present a two-step approach to establish the long context capability of Llama3-70B. First, we extend Llama3-70B base’s context window from 8K to 128K by continually pretraining it on a mix of SlimPajama (Soboleva et al., 2023) with upsampled long sequences (Fu et al., 2024). Then, we apply a three-stage instruction tuning process on curated datasets to enhance the instruction-following, RAG, and long context understanding capabilities at each respective stage. We find that stage-wise instruction-tuning, by incorporating previous datasets, simplifies experimentation and hyperparameter tuning. This approach enhances long context capabilities while maintaining RAG performance.
@@
    We demonstrate that the resulting $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B\text{-}128K}$ can be on par or slightly worse than GPT-4-Turbo-2024-04-09 on many real-world long context understanding tasks. In addition, it outperforms GPT-4-Turbo-2024-04-09 on RAG and conversational QA tasks.
@@
    The current RAG pipeline has limitations that can undermine downstream task accuracy: i) top-k chunk-wise retrieval introduces fragmentation of context; and ii) a small top-k leads to low recall, while a larger k introduces too much irrelevant context to the LLM (e.g., see the analysis in Figure 1 of Yu et al., 2024). We find that the state-of-the-art long-context retriever (Wang et al., 2023c; Lee et al., 2024) can largely alleviate these issues and further improve the RAG-based results for long-context understanding tasks.
@@
    In the comparison between RAG and long-context results, we find that the GPT-4-Turbo level long-context model (including our $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$) outperforms the RAG on 32K benchmarks but still underperforms compared to RAG methods on real-world 128K tasks.
We organize the rest of the paper as follows. We discuss related work in § 2. We introduce the continued pretraining for context window extension in § 3 and the three-stage instruction tuning in § 4. We report results in § 7 and conclude the paper in § 8.
@@ 2.1 Long Context LLM
The trend of extending the context window in LLM starts by Claude with 100K token context (Anthropic, 2023). Although the underlying long context techniques behind proprietary models are unclear, the open LLM and research community has developed many methods to extend the context window of LLMs through continued training or fine-tuning (Kaiokendev, 2023; Nijkamp et al., 2023; Chen et al., 2023a; Tworkowski et al., 2023; Chen et al., 2023b; Peng et al., 2023; Xiong et al., 2023; Fu et al., 2024), especially for open-access LLMs (Touvron et al., 2023a, b) based on rotary position embedding (RoPE) (Su et al., 2024).
@@ 2.2 Retrieval-augmented Generation (RAG)
Retrieval with a standalone retriever (e.g., Karpukhin et al., 2020; Wang et al., 2022; Lin et al., 2023; Lee et al., 2024) is a long-standing solution for handling long texts that cannot fit into the context window of language models. In previous work, various retrieval-augmented language models have been proposed (Nakano et al., 2021; Borgeaud et al., 2022; Wang et al., 2023b, a; Guu et al., 2020; Izacard & Grave, 2021; Izacard et al., 2022; Lewis et al., 2020; Huang et al., 2023; Khandelwal et al., 2019; Liu et al., 2024).
Previous dense-embedding-based retrievers only supported limited context windows (e.g., 512 tokens) (e.g., Karpukhin et al., 2020; Wang et al., 2022; Lin et al., 2023). In top-k chunk-wise retrieval, the short chunk size increases context fragmentation. As a result, extending the context window of retrievers has become popular. For example, Jina Embeddings 2(Günther et al., 2023) and Nomic Embed (Nussbaum et al., 2024) support 8K tokens, while E5-mistral-7B (Wang et al., 2023c) and NV-Embed Lee et al. (2024) support 32K tokens.
3 Extending Context Window to 128K
In this section, we present the method to extend the context window from 8K to 128K for Llama3. We prepare our long context pretraining corpus from the Slimpajama (Soboleva et al., 2023) following Fu et al. (2024). We upsample long-context documents with the hyperparameter set as 0.1 to produce 10 billion tokens with sequence length of 128k. Since Llama3 is pretrained with a much higher RoPE base frequency of 500,000 compared to Llama2, we increased the RoPE base frequency to 150M accordingly. We set the batch size to 32 to fit 4 million tokens in a batch and use a learning rate of 3e-5 to train 2000 steps (8B tokens in total).
@@ 3 Extending Context Window to 128K
4 Instruction-Tuning with Long Context Data
@@
Specifically, we implement three stages of instruction-tuning. For the first two stages, we follow ChatQA 1.5 (Liu et al., 2024), where the model is initially trained on 128k high-quality instruction-following datasets, and then trained on a blend of conversational QA data with provided context. However, these two stages involve relatively short contexts, with a maximum sequence length of only 4K tokens. To enhance our model’s capability to handle very long context sequences up to 128k tokens, we collect a long SFT dataset.
@@
This dataset is collected through two categories: 1) For SFT data sequences less than 32k: We leverage existing long-context datasets from LongAlpaca12k, GPT-4 samples from Open Orca  [^1] https://huggingface.co/datasets/Open-Orca/OpenOrca, and Long Data Collections  [^2] https://huggingface.co/datasets/togethercomputer/Long-Data-Collections. 2) For sequence lengths between 32k and 128k: Since it is challenging to collect such SFT samples, we rely on synthetic datasets. We utilize NarrativeQA, which contains both the ground truth summary and semantically related paragraphs. We assemble all the related paragraphs and randomly insert the ground truth summary to simulate a real long document for its QA pairs. Both the full long SFT dataset and the short SFT dataset from the first two stages are then blended for training. We set the learning rate at 3e-5 and the batch size at 32.
5 Long Context Retriever meets Long Context LLM
@@
As we mentioned in previous section, the current RAG pipeline for LLM has the following issues: i) The top-k chunk-wise retrieval introduces non-negligible fragmentation of context for generating accurate answers. For example, previous state-of-the-art dense-embedding based retrievers (e.g., Li et al., 2023; Lin et al., 2023) only support 512 tokens. ii) Small top-k (e.g., 5 or 10) usually leads to relatively low recall, while much larger k (e.g., 100) can lead to worse generation (see Table 5 in Xu et al. (2024)) as the previous LLMs could not utilize too many chunked context very well (Liu et al., 2023a). To address the issue, we propose to use the most recent long-context retriever (Wang et al., 2023c; Lee et al., 2024), which can support thousands of tokens. In our setting, we use the E5-mistral embedding model (Wang et al., 2023c) as the retriever.
Table 1 compares different chunk sizes for top-k retrieval and the total number of tokens in the context window. Comparing total tokens from 3000 to 12000, we found that more tokens consistently yield better results, confirming the strong long-context capability of our model. We also found that 6000 total tokens offer a good trade-off between cost and performance. With the total number of tokens set to 6000, we discovered that larger chunk sizes give better results. Therefore, we use a chunk size of 1200 and top-5 chunks as the default in our experiments.
chunk-size 300 300 300 600 1200 1200
top-k 10 20 40 10 5 10
total tokens 3000 6000 12000 6000 6000 12000
Avg. 46.31 46.28 46.96 46.88 47.08 47.13
Table 1: Ablation of $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ with RAG given different top-k = {5, 10, 20, 40} retrieval, and chunk-size = {300, 600, 1200} on medium-long context benchmarks within 32K tokens (see Section 6.2 for more details). Given the same budget of total tokens in the context window (e.g., 6000), larger chunk-size (e.g., 1200) gives better results than small chunk-size (e.g., 300 and 600). The accuracy can also improve with larger total tokens in the context window.
6 Evaluation Benchmarks

### v3

@@ Abstract
In this work, we introduce ChatQA 2, an Llama 3.0-based model with a 128K context window, designed to bridge the gap between open-source LLMs and leading proprietary models (e.g., GPT-4-Turbo-2024-04-09) in long context understanding and retrieval-augmented generation (RAG) capabilities. These two capabilities are complementary to each other and essential for LLMs to process large volumes of information that cannot fit into a single prompt. We present a detailed continued training recipe to extend the context window of Llama3-70B-base from 8K to 128K tokens, along with a three-stage instruction tuning process to enhance the model’s instruction-following, RAG performance, and long-context understanding capabilities. Our results demonstrate that the $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ model outperforms most existing state-of-the-art models, including GPT-4-Turbo-2024-04-09, Qwen2-72B-Instruct, and Llama3.1-70B-Instruct, on ultra-long tasks beyond 100K tokens, as well as on the RAG benchmark using only a 4K context window, showing the strong long context capability across varying sequence lengths. We further provide extensive comparisons between direct long-context and RAG solutions using the same state-of-the-art long-context LLMs. Interestingly, we find that the performance of strong long-context LLMs using RAG improves when retrieving a larger number of chunks. With a large set of top-k chunks, RAG consistently outperforms direct long-context solution using the same state-of-the-art long-context models (e.g., $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ and Qwen2-72B-Instruct) on both 32K and 128K benchmarks. We open-source the model weights, training data, and the evaluation setup for the for the community: https://chatqa2-project.github.io/
@@ 1 Introduction
In recent developments, the trend of extending the context window length in LLMs has gained remarkable traction within both the industrial and research communities. All leading proprietary LLMs support very large context window, allowing them to accommodate several hundred pages of text in a single prompt. For example, GPT-4 Turbo (OpenAI, 2023) and Claude 3.5 Sonnet offer a 128K and 200K context window, respectively. Meanwhile, Gemini 1.5 Pro (Gemini-Team, 2024) impressively supports up to a 10M context. Open-access LLMs have also made significant strides to keep up (01.AI et al., 2024; Alibaba-QWen, 2024). For instance, Qwen2-72B-Instruct (Alibaba-QWen, 2024) and Yi-34B (01.AI et al., 2024) support 128K and 200K context windows, respectively. The concurrent work Llama-3.1-70B-Instruct (Meta-AI, 2024) also supports a 128K context window. However, the absence of training data and reproduction recipe makes replicating these models challenging. In addition, these models have mostly been evaluated on synthetic tasks, like Needle in a Haystack (Kamradt, 2023) test, which does not accurately represent real-world downstream task performance. In this work, we aim to create an open source long context recipe that poses the capability at the level of proprietary models (e.g., GPT-4-Turbo-2024-04-09) while focusing on real-world long-context understanding tasks.
Model Type Long Context LLM (128K) Ultra-long (>100K) Long (32K) Short (4K)
4 tasks 6 tasks 9 tasks (RAG)
[Uncaptioned image] GPT-4-Turbo-2024-04-09 33.16 51.93 54.72
[Uncaptioned image] Qwen2-72B-Instruct 39.77 49.94 54.06
Llama3.1-70B-Instruct 39.81 49.92 52.12
[Uncaptioned image] $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ 41.04 48.15 56.30
Table 1: $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ versus state-of-the-art long context LLMs across tasks with varying context lengths. Notably, we open-source the training data and reproduction recipe for building 128K long-context LLMs from pretrained base models with 8K context, resources that are not currently available for open-access (open-weights) 128K context models like Qwen2 and Llama3.1. Our model achieves the highest average score on four real-world ultra-long context (beyond 100K) tasks from InfiniteBench, and nine short context (within 4K) tasks from ChatRAG Bench.
@@
On the other hand, the long-context capability of LLMs is sometimes considered a rival technique to retrieval-augmented generation (RAG). In fact, RAG and long-context techniques complement each other from a pragmatic perspective. An LLM with a long context window can either process large volumes of text as a prompt or utilize retrieval methods to efficiently extract relevant information from the extensive text, depending on the downstream tasks and accuracy vs. efficiency trade-offs. RAG has efficiency advantages and can easily retrieve relevant contexts for query-based tasks (e.g, QA) from billions of tokens, a feat that long context models cannot achieve so far. Meanwhile, long context models are good at tasks such as summarizing entire documents, where RAG can not perform. As a result, the state-of-the-art (SOTA) LLM needs to excel at both capabilities, providing options for different downstream tasks and meeting different accuracy-efficiency trade-off requirements. To compare the solution between long context and RAG, Xu et al. (2024) extended the context window of Llama2 (Touvron et al., 2023b) to 16K and 32K tokens, and studied the interplay between RAG and long-context LLMs. It demonstrates that the RAG method can improve the generation accuracy and inference efficiency of a GPT-3.5-turbo-16k level long context model for QA and query-based summarization tasks.
In this work, we present ChatQA 2, which extends the study of long-context vs RAG by pushing long-context LLMs to the GPT-4-Turbo level capability (with a 128K context window), and combines it with a state-of-the-art long-context retriever for RAG. See Table 1 for an overview of the results. The training data and reproduction recipe for this work are publicly accessible.
@@
    We present a two-step approach to establish the long context capability of Llama3-70B. First, we extend Llama3-70B base’s context window from 8K to 128K by continually pretraining it on a mix of SlimPajama (Soboleva et al., 2023) with upsampled long sequences (Fu et al., 2024). Then, we apply a three-stage instruction tuning process on curated datasets to enhance the instruction-following, RAG, and long context understanding capabilities at each respective stage. We find that stage-wise instruction-tuning, by incorporating previous datasets, simplifies experimentation and hyperparameter tuning.
@@
    We demonstrate that the resulting $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ with 128K context window achieves better accuracy than most existing state-of-the-art models, including GPT-4-Turbo-2024-04-09, Qwen2-72B-Instruct, and Llama3.1-70B-Instruct, on both long-context tasks exceeding 100K tokens and the RAG benchmark for conversational QA tasks within a 4K context window.
@@
    Previous RAG approachs has two potential issues that can undermine downstream task accuracy: i) top-k chunk-wise retrieval introduces fragmentation of context, and ii) a small top-k leads to low recall, while a larger k introduces too much irrelevant context to the LLM (e.g., see the analysis in Figure 1 of Yu et al., 2024). In contrast, we find that the RAG performance of our $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ is highly robust to variations in chunk size when using a long-context retriever (Wang et al., 2023c; Lee et al., 2024). Even more promising, accuracy consistently improves as the total number of retrieved tokens for input increases.
@@
    We conduct extensive comparisons between direct long-context and RAG solutions using the same state-of-the-art long-context LLMs. We find that RAG can still outperform GPT-4-Turbo-level long-context models (including our $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$) on both 32K benchmarks and real-world 128K tasks, given a sufficient number of top-k chunks.
We organize the rest of the paper as follows. We discuss related work in § 2. We introduce the continued pretraining for context window extension in § 3.1 and the three-stage instruction tuning in § 3.2. We report results in § 5 and conclude the paper in § 7.
@@ 2.1 Long Context LLM
The trend of extending the context window in LLM starts by Claude with 100K token context (Anthropic, 2023). Although the underlying long context techniques behind proprietary models are unclear, the open LLM and research community has developed many methods to extend the context window of LLMs through continued training or fine-tuning (Kaiokendev, 2023; Nijkamp et al., 2023; Chen et al., 2023a; Tworkowski et al., 2023; Chen et al., 2023b; Peng et al., 2023; Xiong et al., 2023; Fu et al., 2024), especially for open-access LLMs (Touvron et al., 2023a; b) based on rotary position embedding (RoPE) (Su et al., 2024).
@@
Note that, the Llama-3.1 models (Dubey et al., 2024) with a 128K context window were introduced after the release of this work. We add a comprehensive comparison between ChatQA-2 and Llama-3.1 in the experiments.
@@ 2.2 Retrieval-augmented Generation (RAG)
Retrieval with a standalone retriever (e.g., Karpukhin et al., 2020; Wang et al., 2022; Lin et al., 2023; Lee et al., 2024) is a long-standing solution for handling long texts that cannot fit into the context window of language models. In previous work, various retrieval-augmented language models have been proposed (Nakano et al., 2021; Borgeaud et al., 2022; Wang et al., 2023b; a; Guu et al., 2020; Izacard & Grave, 2021; Izacard et al., 2022; Lewis et al., 2020; Huang et al., 2023; Khandelwal et al., 2019; Liu et al., 2024).
In general, the input to a retriever or dense embedding model is a chunk of text (i.e., a sequence of subword tokens), and the output is an embedding vector. These dense embedding vectors are then indexed and retrieved using k-nearest-neighbor retrieval. Previous dense-embedding-based retrievers only supported limited chunk size (e.g., 512 tokens) (e.g., Karpukhin et al., 2020; Wang et al., 2022; Lin et al., 2023) due to the context window size of pretrained BERT base model. In top-k chunk-wise retrieval, the short chunk size increases context fragmentation. As a result, extending the context window of retrievers has become popular. For example, Jina Embeddings 2(Günther et al., 2023) and Nomic Embed (Nussbaum et al., 2024) support 8K tokens, while E5-mistral-7B (Wang et al., 2023c) and NV-Embed (Lee et al., 2024) support 32K tokens.
3 Method
3.1 Extending Context Window to 128K
In this section, we present the method to extend the context window from 8K to 128K for Llama3. Note that, the Llama-3.1 models with 128K context window are released after this work. We compare our $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ and $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}8B}$ with Llama-3.1-70B-Instruct and Llama-3.1-8B-Instruct in experiments.
We prepare our long context pretraining corpus from the Slimpajama (Soboleva et al., 2023) following Fu et al. (2024). We upsample long-context documents with the hyperparameter set as 0.1 to produce 10 billion tokens with sequence length of 128k. Since Llama3 is pretrained with a much higher RoPE base frequency of 500,000 compared to Llama2, we increased the RoPE base frequency to 150M accordingly. We set the batch size to 32 to fit 4 million tokens in a batch and use a learning rate of 3e-5 to train 2000 steps (8B tokens in total).
@@ 3.1 Extending Context Window to 128K
3.2 Instruction-Tuning with Long Context Data
@@
Specifically, we implement three stages of instruction-tuning. For the first two stages, we follow ChatQA 1.5 (Liu et al., 2024), where the model is initially trained on a high-quality instruction-following datasets, and then trained on a blend of conversational QA data with provided context. However, these two stages involve relatively short contexts, with a maximum sequence length of only 4K tokens. To enhance our model’s capability to handle very long context sequences up to 128k tokens, we collect a long SFT dataset.
@@
This dataset is collected through two categories: 1) For SFT data sequences less than 32k: We leverage existing long-context datasets from LongAlpaca12k, GPT-4 samples from Open Orca  [^1] https://huggingface.co/datasets/Open-Orca/OpenOrca, and Long Data Collections  [^2] https://huggingface.co/datasets/togethercomputer/Long-Data-Collections. 2) For sequence lengths between 32k and 128k: Since it is challenging to collect such SFT samples, we rely on synthetic datasets. We utilize NarrativeQA (Kočiskỳ et al., 2018) [^3] https://huggingface.co/datasets/deepmind/narrativeqa which includes summary paragraphs, questions, answers, and source long web pages. The summaries are human-generated based on the source web pages, while the question-answer pairs are annotated by humans using the summaries. To extend the context length, we inserted a summary into the corresponding long web page document at a random location, ensuring that sentence structure was not disrupted. This approach maintains the grounding of the question-answer pairs within the augmented long documents.. Since we used NarrativeQA for synthetic data generation, we intentionally excluded it from the evaluation benchmarks to avoid any potential data contamination. Both the full long SFT dataset and the short SFT dataset from the first two stages are then blended for training. We set the learning rate at 3e-5 and the batch size at 32.
3.3 Long Context Retriever meets Long Context LLM
@@
As we mentioned in previous section, the current RAG pipeline for LLM has the following issues: i) The top-k chunk-wise retrieval introduces non-negligible fragmentation of context for generating accurate answers. For example, previous state-of-the-art dense-embedding based retrievers (e.g., Li et al., 2023; Lin et al., 2023) only support 512 tokens. ii) Small top-k (e.g., 5 or 10) usually leads to relatively low recall, while much larger k (e.g., 100) can lead to worse generation (see Table 5 in Xu et al. (2024)) as the previous LLMs could not utilize too many chunked context very well (Liu et al., 2023a). To address the issue, we propose to use the most recent long-context retriever (Wang et al., 2023c; Lee et al., 2024), which can support thousands of tokens. In our setting, we use the E5-mistral embedding model (Wang et al., 2023c) as the retriever. The input of E5-Mistral model is a chunk of text represented by a sequence of subword tokens, and the output is an embedding vector. The long documents or corpus are chunked and embedded into a set of dense embedding vectors. Such dense embedding vectors are indexed and retrieved by k-nearest-neighbor retrieval at inference time, where the embedding of prompt / question is treated as query.
4 Baselines and Evaluation Benchmarks
Long context models